OpenStudy (anonymous):

Find the function f(x) which has the following Laplace transform: F (s) = (4s+3)/(s^2-2s+8)

OpenStudy (anonymous):

lol this is bringing back my differential equations days...i'll try

OpenStudy (anonymous):

You have to take the inverse laplace transform of F(s)

OpenStudy (anonymous):

haha. Personally I find it a hateful section! :) Don't you have to break up the fraction in some way. If the 8 on the bottom was negative it would be so much easir!! :)

OpenStudy (anonymous):

(partial fraction) but i cant seem to do it...

OpenStudy (anonymous):

Yeah the bottom function is not factorable so we'll need to complete that expression as a complete square

OpenStudy (anonymous):

Okay when you complete the square in the denominator you get: s^2 - 2x + 8 = (s - 1)^2 + 7

OpenStudy (anonymous):

I'm sure that the laplace will transform into either sin or cos so we're going to work on the top as well

OpenStudy (anonymous):

[4(s - 1) + 4 + 3 ] / (s - 1)^2 + 7

OpenStudy (anonymous):

4(s - 1) / (s - 1)^2 +7 Remember that 7 = [sqrt(7)]^2

OpenStudy (anonymous):

Also we have 7 / [(s-1)^2 + 7]

OpenStudy (anonymous):

For 4(s-1) / (s -1)^2 + 7 the inverse laplace transform is \[4e ^{t}\cos (\sqrt{7}t)\]

OpenStudy (anonymous):

For the second one: 7 = sqrt[7] x sqrt [7]

OpenStudy (anonymous):

So we can write it as \[\sqrt{7}[\sqrt{7}/[(s - 1)^2 + (\sqrt{7})^{2}]\]

OpenStudy (anonymous):

So that leaves us with \[\sqrt{7}e ^{t}\sin(\sqrt{7}t)\]

OpenStudy (anonymous):

So your overall answer is f(t) = 4e^tcos(7√t) + √7e^tsin(7√t) Do you get it?

OpenStudy (anonymous):

YES!!! I DID!!!! I can't thank you enough!!! Wish I could give you more than one medal!!!! :)

OpenStudy (anonymous):

No problem! It's good you understand...I thought I will have to attach a file because explanation seemed long